# Cortex AISQL: A Production SQL Engine for Unstructured Data

### v2

@@ 7. Related Work
The intersection of databases and AI has evolved along two directions: AI4DB, which utilizes ML to enhance database systems themselves, and DB4AI, which integrates native AI capabilities into databases to simplify analytics and model inference. AISQL builds upon both paradigms by integrating LLMs directly into the SQL execution layer.
@@
The AI4DB line of research applies ML algorithms to internal database components such as query optimization, indexing, and configuration tuning. Early systems like LEO (stillger2001leo) used feedback learning to refine cost estimates for the DB2 optimizer. Later, OtterTune (vanaken2017ottertune) and CDBTune (zhang2019end) leveraged large-scale supervised and reinforcement learning to automatically tune database parameters, inspiring the vision of self-driving databases (pavlo2017self).
@@
A rich body of work explores learned query processing components. Learned Indexes (kraska2018case) replaced traditional index structures with neural networks, while ALEX (ding2020alex) extended these ideas to support dynamic workloads. In query optimization, learned cardinality estimation models, such as MSCN (kipf2018learned), Naru (yang2019deep), and DeepDB (hilprecht2020deepdb), achieved notable improvements in selectivity prediction by learning correlations across attributes and tables. Reinforcement learning has also been applied to plan search and join ordering, exemplified by ReJOIN (marcus2018deep), Neo (marcus2019neo), and Bao (marcus2021bao). These works culminate in the SageDB (kraska2019sagedb), which proposes a learned database system where each component is automatically specialized for a given data distribution and workload.
@@
The application of ML to query optimization extends to handling expensive predicates in specialized domains. Systems such as BlazeIt (kang2019blazeit) and TASTI (kang2020tasti) optimize queries involving expensive ML-based predicates using model cascades, semantic indexes, and query-specific proxy models. In particular, model cascades that combine cheap approximate models with expensive accurate ones (kang2020approximate; chen2023frugal; zellinger2025rational) have influenced the design of AISQL’s approach to managing expensive AI operations.
The DB4AI perspective takes the inverse view: enabling databases to host and execute ML models or reasoning workloads natively, rather than treating them as external applications. This line of work seeks to unify analytics and model inference under the relational abstraction, with the goal of allowing users to express AI tasks declaratively using SQL.
@@
Earlier in-database ML frameworks such as MADLib (hellerstein2012madlib) extended SQL with support for descriptive statistics and simple ML methods. Recent systems go further by integrating semantic and language-model capabilities directly into query processing, for example LOTUS (patel2025semantic) introduces a set of semantic operators for Pandas-like dataframe processing. ThalamusDB (jo2025thalamusdb) explores approximate query processing for multimodal data, supporting natural language predicates over visual, audio, and textual content through zero-shot models combined with relational operators.
@@
Several systems have proposed specialized approaches to LLM-powered SQL processing. SUQL (liu2024suql) augments SQL with answer and summary operators for knowledge-grounded conversational agents, with the focus on row-wise LLM operations for question answering. ZenDB (lin2024towards) optimizes SQL queries for extracting structured data from semi-structured documents with predictable templates. Palimpzest (liu2025palimpzest) provides a declarative framework for LLM-powered data processing with specialized map-like operations and basic cost-based optimizations. UQE (dai2024uqe) studies embedding-based approximations for LLM-powered filters and stratified sampling for aggregations, though it provides best-effort performance without accuracy guarantees.
@@
Beyond SQL-centric approaches, several frameworks have emerged for LLM-powered data processing. Aryn (anderson2024design) offers a Spark-like API with PDF extraction and human-in-the-loop processing capabilities. DocETL (shankar2025docetl) proposes agent-driven pipeline optimization for complex document processing tasks and uses LLM agents to explore task decomposition strategies. EVAPORATE (arora2023language) specializes in extracting semi-structured tables through code synthesis. These systems demonstrate the potential of LLM-powered processing, but often lack formal cost models and provide limited performance guarantees.
Optimizing expensive predicates and user defined functions was the subject of a number of research papers in the past (hellerstein6caching) (hellestein93predicate) (chaudhuri99optimization). This line of work inspired many of the optimizations performed in AISQL. In (hellestein93predicate) a cost-based optimization technique was introduced that improved query performance by selectively reordering or “migrating” expensive selection predicates in relational query plans. The authors developed an analytical model and algorithm that balances predicate selectivity and evaluation cost, achieving faster query execution without exhaustive reordering. (hellerstein6caching) presents strategies to optimize query performance by caching the results of costly user-defined or external function calls within a query. It introduces algorithms that determine when and how to reuse previously computed method results, reducing redundant evaluations and significantly improving overall query efficiency. Finally, (chaudhuri99optimization) proposed techniques to integrate expensive, user-defined predicates into traditional cost-based query optimization frameworks. It developed a formal model for estimating predicate costs and selectivities, enabling the optimizer to reorder and combine such predicates effectively to minimize overall query execution time.
AISQL unifies the goals of both research directions. From AI4DB, it borrows the principle of learning to optimize workloads by applying cost modeling, feedback, and adaptive execution to manage expensive semantic operators. From DB4AI, it inherits the vision of integrating LLM-powered inference directly into the SQL engine. Unlike systems that provide best-effort performance or focus solely on row-wise operations, AISQL co-optimizes query execution and LLM inference within a unified relational framework. AISQL represents a step toward production-ready AI-native database systems that reason, learn, and optimize within the same execution environment.
@@ 8. Conclusions
The proliferation of unstructured data alongside traditional structured datasets has created a fundamental gap in modern data warehouses: users need to perform semantic reasoning over hybrid data but are forced to export data to external systems, orchestrate complex pipelines, and sacrifice the optimization benefits of integrated query processing. This paper presents Snowflake AISQL, a production SQL engine that bridges this gap by integrating LLM capabilities directly into SQL through semantic operators.
AISQL introduces six semantic operators that compose naturally with relational SQL primitives. These operators enable users to express complex analytical tasks that blend structured queries with semantic reasoning in a single declarative statement, and eliminate the need for external data movement and manual pipeline orchestration.
The core technical challenge we address is making semantic operators efficient at a production scale. LLM-based semantic operators are orders of magnitude more expensive than traditional SQL operations; their cost and selectivity are opaque to the query optimizer, and conventional optimization heuristics can produce plans that are catastrophically inefficient. Our solution comprises three techniques driven by analysis of real customer workloads. AI-aware query optimization treats LLM inference cost as a first-class objective, resulting in efficient operator placement and predicate evaluation order. Adaptive model cascades reduce costs through intelligent routing between lightweight proxy models and powerful oracle models. Query rewriting for semantic joins transforms quadratic-complexity joins into a linear multi-label classification task and, as a result, delivers speedups with improved prediction quality.
Our production deployment at Snowflake validates these design choices. Analysis of customer workloads reveals that AI operators consume the majority of execution costs, and nearly half of queries involve multi-table operations. These observations directly motivate our focus on join optimization and cost-aware planning. Extensive experimental evaluation across diverse benchmarks demonstrates that AISQL successfully delivers both efficiency and quality at the production scale.
AISQL represents a step toward unified data processing architectures where semantic reasoning over unstructured content is a first-class database capability rather than an external add-on. By integrating LLM capabilities into the SQL engine, with a focus on cost optimization, we enable organizations to unlock insights from their data assets through familiar declarative interfaces, while maintaining the performance, scalability, and reliability expected from modern data warehouses.

### v3

@@ 7. Related Work
The AI4DB line of research applies ML algorithms to internal database components such as query optimization, indexing, and configuration tuning. Early systems like LEO (Stillger et al., 2001) used feedback learning to refine cost estimates for the DB2 optimizer. Later, OtterTune (Van Aken et al., 2017) and CDBTune (Zhang et al., 2019) used large-scale supervised and reinforcement learning to automatically tune database parameters, inspiring the vision of self-driving databases (Pavlo et al., 2017).
@@
A rich body of work explores learned query processing components. Learned Indexes (Kraska et al., 2018) replaced traditional index structures with neural networks, while ALEX (Ding et al., 2020) extended these ideas to support dynamic workloads. In query optimization, learned cardinality estimation models, such as MSCN (Kipf et al., 2018), Naru (Yang et al., 2019), and DeepDB (Hilprecht et al., 2020), achieved notable improvements in selectivity prediction by learning correlations across attributes and tables. Reinforcement learning has also been applied to plan search and join ordering, exemplified by ReJOIN (Marcus and Papaemmanouil, 2018), Neo (Marcus et al., 2019), and Bao (Marcus et al., 2022). These works culminate in the SageDB (Kraska et al., 2019), which proposes a learned database system where each component is automatically specialized for a given data distribution and workload.
@@
The application of ML to query optimization extends to handling expensive predicates in specialized domains. Systems such as BlazeIt (Kang et al., 2019) and TASTI (Kang et al., 2022) optimize queries involving expensive ML-based predicates using model cascades, semantic indexes, and query-specific proxy models. Model cascades that combine cheap approximate models with expensive accurate ones (Kang et al., 2020; Chen et al., 2023; Zellinger and Thomson, 2025) have influenced the design of AISQL’s approach to managing expensive AI operations.
@@
Earlier in-database ML frameworks such as MADLib (Hellerstein et al., 2012) extended SQL with support for descriptive statistics and simple ML methods. Recent systems go further by integrating semantic and language-model capabilities directly into query processing. LOTUS (Patel et al., 2025) introduces a set of semantic operators for Pandas-like dataframe processing. ThalamusDB (Jo and Trummer, 2024) explores approximate query processing for multimodal data, supporting natural language predicates over visual, audio, and textual content through zero-shot models combined with relational operators. Unlike these systems, AISQL operates within a production distributed SQL engine and co-optimizes query planning with LLM inference cost.
@@
Several systems have proposed specialized approaches to LLM-powered SQL processing. SUQL (Liu et al., 2024) augments SQL with answer and summary operators for knowledge-grounded conversational agents, with the focus on row-wise LLM operations for question answering. ZenDB (Lin et al., 2024) optimizes SQL queries for extracting structured data from semi-structured documents with predictable templates. Palimpzest (Liu et al., 2025) provides a declarative framework for LLM-powered data processing with specialized map-like operations and basic cost-based optimizations. UQE (Dai et al., 2025) studies embedding-based approximations for LLM-powered filters and stratified sampling for aggregations, though it provides best-effort performance without accuracy guarantees. AISQL goes beyond row-level processing with cross-row optimizations (cascades, join rewriting) that reduce inference calls by orders of magnitude.
@@
Beyond SQL-centric approaches, several frameworks have emerged for LLM-powered data processing. Aryn (Anderson et al., 2024) offers a Spark-like API with PDF extraction and human-in-the-loop processing capabilities. DocETL (Shankar et al., 2025) proposes agent-driven pipeline optimization for complex document processing tasks and uses LLM agents to explore task decomposition strategies. EVAPORATE (Arora et al., 2023) specializes in extracting semi-structured tables through code synthesis. These systems demonstrate the potential of LLM-powered processing, but often lack formal cost models and provide limited performance guarantees.
Optimizing expensive predicates and user-defined functions has been extensively studied (Hellerstein and Naughton, 1996; Hellerstein and Stonebraker, 1993; Chaudhuri and Shim, 1999) and directly inspired AISQL’s query optimizations. Prior work addressed cost-based predicate reordering (Hellerstein and Stonebraker, 1993), caching of expensive function results (Hellerstein and Naughton, 1996), and integrating user-defined predicates into cost-based optimization frameworks (Chaudhuri and Shim, 1999). AISQL extends these techniques to LLM-powered operators, where per-row costs are orders of magnitude higher and selectivities are unknown at compile time.
@@ 8. Conclusions
We have presented Snowflake AISQL, a production SQL engine that bridges the gap between structured and unstructured data processing by integrating LLM capabilities directly into SQL. AISQL introduces six semantic operators that compose naturally with relational SQL primitives, enabling users to express complex analytical tasks that blend structured queries with semantic reasoning in a single declarative statement.
The core technical challenge we address is making semantic operators efficient at production scale. Our solution comprises three techniques driven by analysis of real customer workloads: AI-aware query optimization that treats LLM inference cost as a first-class objective; adaptive model cascades that reduce costs through intelligent routing between lightweight proxy and powerful oracle models; and query rewriting for semantic joins that transforms quadratic-complexity joins into linear multi-label classification tasks with improved prediction quality.
Our production deployment at Snowflake validates these design choices, and experimental evaluation demonstrates that AISQL delivers both efficiency and quality at production scale. AISQL represents a step toward unified data processing architectures where semantic reasoning over unstructured content is a first-class database capability. Organizations can query both structured and unstructured data through familiar declarative interfaces while maintaining the performance and reliability expected from modern data warehouses.
Several directions for future work remain. Caching runtime statistics across queries would improve plan quality for recurring AISQL workloads. Extending model cascades beyond AI_FILTER to multi-class operators requires generalizing the binary threshold framework to handle distinct confidence distributions per class. Hybrid join strategies that combine classification-based rewriting with filtering could improve recall on datasets where the rewrite alone sacrifices coverage.